INAF queue observing, 2018 Jan 22-23 
We (the observers) take this space to bring attention of the world to the the awkward situation of large part
of research personnel in INAF.
We are now performing observations for INAF researchers with LBT, one of the biggest telescope in the world.
But we are also valued for our scientific contribution, for results that we publish as first authors and for those in which we are involved.
However, we believe that like for many other INAF post-docs,  
our contract situation and our expectations do not reflect the great results that we, and INAF thanks to us, are obtaing.
Like many other post-docs, we are paid on contracts renewed every  year
pending the availability of external funds.
Moreover, we are not entitled to officially say a word on institutional choice and in many aspects
we are not considered as INAF employees, but like external collaborators.
Indeed, we cannot even lead INAF projects and build a independent research.
Now, we believe INAF has given the power and instruments to start changing this situation.
Observer: R. Carini, A. Rossi 
Telescope Operator:  S. Allanson  
Support Astronomer: B. Rothberg  on call 
Instrument:  LBC
  Summary 
Night shared with LINC team.
Instruments: LBC works, LUCI1 only imaging, LUCI2 imaging, LS and MOS, AO works just for LUCI1 but the camera N30 has problems, MODS1  works in both channels, MODS2 works just in the blue channel.
Night duration (12\xB0 el twilight)  2:00am- 13:15pm UT = 
11 h 15m
Night shared with LINC team. INAF time  7:00pm- 13:15pm UT = 
6 h 15m , The LNC team finishes early, we start at 7:00 pm.  
Observing time: 
6 h   (96%)
Weather time loss: 
0 h  (0%)
technical time loss 
15 m (4%)
  Data Summary 
	
		
			|  Project_RUN  | 
			 Objects  | 
			 Instrument  | 
			 Filter  | 
			 ExpTime  | 
			 Notes  | 
		
	
	
		
			|  Annibali R13  | 
			 UGC 4459  | 
			 LBC  | 
			 g,r  | 
			 1200s x 3 in both g,r  | 
			 completed  | 
		
		
			|  Annibali R15  | 
			 NGC2976  | 
			 LBC  | 
			 g,r  | 
			 1200s x 3 in both g,r  | 
			 completed  | 
		
		
			|  Palazzi_29  | 
			 GW170817  | 
			 LBC  | 
			 r,r  | 
			 1h  | 
			 completed  | 
		
		
			|  Sky Flat  | 
			    | 
			 LBC  | 
			 r,r  | 
			    | 
			    | 
		
		
			|  Bias  | 
			    | 
			 LBC  | 
			    | 
			    | 
		
	
  Details 
07:00 slew to Annibali R13 w seeing 0.7 
-  dohybrid
  -  07:16 copointing
  -  07:20 calibration ob UGC4459_calib
  -  seeing 0.7 on dimm
  -  07:30 ob UGC4459_1
  -  seeing 0.7 on image at 7:35 UT
  -  
Roberta says: il rosso e' molto bello
  -  seeing 0.7 on image at 7:50 UT
  -  dofpia
 
 
07:58 ob UGC4459_2 
-  08:09 seeing on image 0.68''
 
 
08:26 ob UGC4459_3 
-  08:38 seeing on image 0.8"
  -  08:50 seeing on image 0.8" but last exposure stars have triangle shape
  -  dofpia
  -  run again to get other 4 exposures with good collimation
 
 
09:09 slew to R15 
-  dofpia
  -  copinting
  -  dimm seeing 0.78"
  -  09:21* NGC2976_calib.ob
  -  09:25 seeing on image 0.8"
 
 
09:26 NGC2976_1.ob 
-  09:33 seeing on image 0.9"
  -  dimm seeing 0.75"
  -  dofpia
 
 
09:58 NGC2976_2.ob 
-  10:05 seeing on image between 0.8" and 1.0''
  -  10:24 seeing on image 0.8"
  -  dofpia
 
 
10:33 NGC2976_3.ob 
-  10:45 seeing on image 0.8"
  -  10:57 seeing on image 0.8"
  -  
105655 lbc red image has readout problem, probably not useful!
  -  we take another exposure
 
 
11:08 Slew to Palazzi_29 GW170817 
-  the screen frozen, we change station, we pass to obs4.
 
 
11:16 slew to Palazzi_29 target GW170817 
-  dofpia, the target is very low, the field is not good to find the convergence, the TO makes an offset and we try again the dofpia.
  -  11:27 dofpia works
  -  11:31 run copointing
  -  11:35 lbcrangebal crash
  -  11:40 we come back to obs2
  -  11:45 we send lbcrangebal again, now it works
 
 
11:47 run ob LBT-2017B-C0831_2science.xml 
-  seeing on image 0.7''
  -  115527 lbc blue image with bad readout column, but non over the target
  -  12:00 seeing on image 0.8''
  -  12:07 stars are elongated, stop e run dofpia
  -  12:16 run again the ob LBT-2017B-C0831_2science.xml
  -  12:28 seeing 0.6" on image
 
 
12:54 dofpia 
-  12:57 run the ob LBT-2017B-C0831_2science.xml
  -  01:06 seeing on image 0.6''-0.7''
  -  good background until lbcr. 20180123.130932
  -  01:30 stop the integration
 
 
01:40 start with the  sky flat,  good r in blue and r in red
02:10 do25bias
-- %USERSIG{RobertaCarini - 2018-01-24}%